News by Date
News by Category
Solar System
Stars
White Dwarfs
Supernovas
Neutron Stars
Black Holes
Milky Way Galaxy
Normal Galaxies
Quasars
Galaxy Clusters
Cosmology/Deep Field
Miscellaneous
Press Resources
Status Reports
Press Advisories
Image Releases
Release Guidelines
Image Use Policy
NASA TV
Biographies/Interviews
Web Shortcuts
Chandra Blog
RSS Feed
Chandra Mobile
Chronicle
Email Newsletter
News & Noteworthy
Image Use Policy
Questions & Answers
Glossary of Terms
Download Guide
Get Adobe Reader
Chandra @ NASA
Visit the Chandra pages at the NASA portal
Image Use
Image Use Policy & Request Form
Guidelines for utilizing images, applets, movies, and animations featured in this Web Site.
Getting Hard Copies of Images
Ways to obtain photos, slides, etc of Chandra images.
Operations CXO Status Report

Friday 6 April 2001 6.30pm EST

Last week the observing schedule was interrupted following a high radiation event at 10:22pm EDT on Apr 2 that stopped the daily load and activated the science instrument safing sequence (SCS 107). The radiation event followed the largest solar flare in 25 years.

All actions were nominal and assessment of the dump data confirmed that the trigger was due to high proton counts in the EPHIN E1300 channel exceeding the pre-defined threshold of 10 counts/cm^2/s/sr. The solar levels subsided to an acceptable level over the next 72 hours and the loads were replanned (3 times) and resumed at 1.17am EDT on 6 Apr. The observations of 1928+738, 1641+399, the XMM 13 hr Survey Field and V-1121+2327 were impacted and will be rescheduled in a future load. The load was resumed in time for a coordinated observation of the Crab Nebula with the Hubble Space Telescope.

A real-time procedure was completed on Apr 6 to insert the LETG grating. Telemetry indicated a nominal insertion, and the retraction is scheduled for Apr 7.

The planned observations for the coming week are shown below and include an observation of NGC 3516 with XMM, RXTE, and BeppoSAX, and an observation of Mkn 509 with the Hubble Space Telescope.

------------------------------------------
ISO NW #2               ACIS-I       Apr 8  
NGC 3516                ACIS-S/HETG
Radiation Belts                      Apr 9  
NGC 3516                ACIS-S/HETG  Apr 10 
Radiation Belts                      Apr 12 
Mkn 509                 ACIS-S/HETG  Apr 13 
EXO 0748-676            ACIS-S/HETG
A3558                   ACIS-S       Apr 14 
NGC 5018                ACIS-S
TWA-5                   ACIS-S       Apr 15 
Radiation Belts

------------------------------------------

All spacecraft subsystems continued to operate nominally.

Return to Status Reports